OpenStudy (amenah8):

FIND FAETA: cos faeta = (-15/17), sin faeta > 0

OpenStudy (michele_laino):

do you mean cos(\theta) and sin(\theta), right?

OpenStudy (amenah8):

Right, sorry!

OpenStudy (michele_laino):

the requested angle \theta lies in the second quadrant, since we have this drawing: 

OpenStudy (michele_laino):